Question 4 Convert the given polar coordinates to rectangular coordinates. (a) (6,35) *)(-2,- ) (c) (1,0)
Answer #1

Solution;

a). The polar point is

=(r,\theta)=(6,-\frac{3\pi}{4})

The rectangular equation is

x= r.cos\theta

y= r.sin\theta

Then

x= 6*cos(-\frac{3\pi}{4})=6*(-\frac{1}{\sqrt{2}})=-3\sqrt{2}

y= 6*sin(-\frac{3\pi}{4})=6*(-\frac{1}{\sqrt{2}})=-3\sqrt{2}

The rectangular equation is

=(-3\sqrt{2},-3\sqrt{2})

b) the polar equation is

=(r,\theta)=(-2,-\frac{3\pi}{2})

Then,

The rectangular equation

x= (-2)cos(-\frac{3\pi}{2})=0

x= (-2)sin(-\frac{3\pi}{2})=-2

Answer =(0,-2)

C) the polar equation is

=(r,\theta)=(1,0)

The rectangular equation is

x= (1)cos(0)=1

And

y= (1)sin(0)=0

Answer (1,0)